# Dense Feature Interaction Network for Image Inpainting Localization

### v1

@@ III-B Reverse Edge Attention Enhancement
Filling content to a specific area in an image can leave detectable traces between the altered region and its surroundings. Therefore, leveraging these edge artifacts is crucial for detecting inpainting, which the mid-lowe-level features can guide. Nevertheless, mid-low-level features not only contain inconsistencies caused by inpainting but also a large amount of detail and structural noise, such as texture, shape, and object boundaries. Using traditional edge extraction operators [14, 53] or convolutional structures [54] will weaken the representation of inpainting artifacts. Leveraging context-rich features that include the structural information of the restoration region as a prior can effectively suppress irrelevant semantic noise and guide boundary learning.
We develop a method for detecting and enhancing edge artifacts using reverse attention called REAE, which uses the structural details in the mid-level feature $f_{3}^{3}$ as a guide, directing the attention of low-level features toward subtle boundary changes of the inpainting area. This promotes learning distinctive features to differentiate between inpainting edges and natural boundaries. As shown in Fig. 3(a), the REAE module accepts mid-level features $f_{3}^{3}$ and low-level features $f_{1}^{3}$ and $f_{2}^{3}$ as input. Its upper branch aggregates structural information of suspicious regions through average pooling, max pooling, and $7\times 7$ convolution. Subsequently, the sigmoid output is inverted to obtain a reverse attention map emphasizing edges, expressed by Eq. (III-B). Meanwhile, the lower branch of the REAE progressively combines $f_{3}^{3}$, $f_{2}^{3}$, and $f_{1}^{3}$ for $f_{123}^{e}$ through $3\times 3$ convolution, Batch Normalization (BN), and bilinear interpolation upsampling. Finally, we use Eq. (3) to fuse the feature for edge supervision and output the edge mask $M_{e}$ through a prediction head composed of multiple convolutional layers.
@@
$Me=Pred_{edge}(f_{123}^{e}\otimes Att_{re}),$ (3)
@@ III-C Spatial Adaptive Feature Fusion
Due to the nature of neural networks, deep features exhibit a significantly stronger response to the target region compared to shallow features.The difference in receptive field inevitably leads the model to bias towards the side with a more robust response during the fusion process, allocating more weights to the deep features, which can severely suppress the role of shallow features. Shallow features can provide detailed edge positioning and structural information, significantly impacting the detection accuracy.
@@
To unify the feature representation of the inpainting region, we design a spatial adaptive feature fusion module (SA-FF) to fuse high-level features and mid-low-level features to correct the bias brought by response intensity. Normalization can limit the response of different features to a similar magnitude, effectively eliminating the interference of response intensity. After normalization, the model can treat deep and shallow features fairly, making it easier to learn effective fusion weights. Specifically, we reduce the channels of $f_{123}$ and $f_{4}$ to 2 through a $1\times 1$ convolution, followed by normalization. The process can be expressed as: Eq. (4) and (5).
$F_{123}^{Norm}=BN(Conv_{1}(f_{123})),F_{123}^{Norm}\in\mathbb{R}^{H\times W\times 2}$ (4)
$F_{4}^{Norm}=BN(Conv_{1}(f_{4})),F_{4}^{Norm}\in\mathbb{R}^{H\times W\times 2}$ (5)
@@
For different spatial locations, there exists a disparity in the contribution of deep and shallow features. Deep features, rich in context, are semantically accurate but lose many spatial and geometric details. Shallow features better retain these aspects, and we hope that the model trusts shallow features more in the boundary area and uses the information provided by deep features to fill the target area. As shown in the Fig. 5, we designed a spatially adaptive weight learner to adjust the fusion weights according to positional features, fully utilizing both advantages. Specifically, we feed the features $f_{w}$ used for learning weights into alternating connection blocks composed of $1\times 1$ convolution, BN, and ReLU to generate position-aware fusion weights $\textbf{W}\in\mathbb{R}^{H\times W\times 4}$. Then, the final fusion output is obtained according to element-wise multiplication and element-wise addition, as shown in Eq. (8). Finally, a region mask $M_{r}$ is output through a prediction head composed of convolutional layers.
$\textbf{W}=((\omega_{1})_{i,j},(\omega_{2})_{i,j},(\omega_{3})_{i,j},(\omega_{4})_{i,j}),i\in[1,H],j\in[1,W]$ (6)
$F^{Norm}=ReCat(F_{123}^{Norm},F_{4}^{Norm}),F^{Norm}\in\mathbb{R}^{H\times W\times 4}$ (7)
$F^{fuse}$ $=Cat(\oplus(F^{Norm}[0:2]\otimes\textbf{W}[0:2]),$
$\oplus(F^{Norm}[2:4]\otimes\textbf{W}[2:4]))$ (8)
@@
where $ReCat(\cdot)$ denotes the concatenation of $F_{123}^{Norm}$ and $F_{4}^{Norm}$ after rearranging them in channel order. $Cat(\cdot)$ represents the ordinary concatenation along the channel dimension. $\otimes$ and $\oplus$ denote the element-wise multiplication and element-wise addition, respectively. []delimited-[][$\cdot$][ ⋅ ] is defined as the slicing operation of the channel dimension. Note that in our design, W will learn different weights for each spatial position of each channel, totaling $4\times H\times W$ weights.
@@ III-D Loss Function
To train our DeFI-Net, we use an objective function that combines the region loss and the edge loss functions. The region loss enhances the model’s sensitivity to the inpainting region from a global perspective. It promotes the overlap between the predicted mask and the ground truth. In contrast, the edge loss guides the model to learn non-semantic detail differences from a local perspective, achieving more precise localization. According to [58], we define the region loss function as $L_{region}=L_{bce}^{w}+L_{iou}^{w}$. Compared with the traditional BCE loss, $L_{bce}^{w}$ assigns different weights to pixels in different positions, paying more attention to the target area and hard pixels in the area (locations such as boundaries prone to classification errors). The IoU loss $L_{iou}^{w}$ assigns greater weight to the target area, prompting the network to pay attention to the global structure. For the edge loss, we use the $L_{bce}$ similar to supervise the edge prediction, denoted as $L_{edge}$. The overall loss function is a convex combination of two losses: $L_{total}=\lambda_{r}L_{region}+\lambda_{e}L_{edge}$.
@@ IV-A Experimental Settings
Datasets. To evaluate the performance of our DeFI-Net, we use five diverse image inpainting datasets, considering variations in inpainting model, region size, and image content. Table I summarizes the datasets. IID-84K [18] comprises 74K inpainted images created using the Gated Convolution (GC) model [59] on images sourced from Places [60] (with JPEG lossy compression) and Dresden [61] (with NEF lossless compression). Additionally, it includes a subset with 1K GC images for in-domain testing (denoted as IID-GC) and another 9K images (denoted as IID-9K) generated by four other deep learning inpainting models (CA [29], SH [28], LB [30], RN [62]) and five traditional inpainting models (TE [63], NS [2], PM [64], SG [65], LR [66]) on images from CelebA [67] and ImageNet [68]. DEFACTO [69] contains different manipulation types while our primary focus lies on 20K inpainting images and 4K authentic images (DE-24K) for training, along with a validation set (DE-1K) and a test set (DE-3K). In addition, we consider the diverse inpainted images in Korus [70], Photoshop [20], and AutoSplice [38] to evaluate the generalization ability of our method.
TABLE I: Details of image inpainting datasets in our experiment.
@@
Implementation Details. We use the HRNet [51] initialized with pre-trained weights from ImageNet[68] as an example. Our model is implemented in Pytorch and trained using the Adam optimizer. During the training phase, we set the size of the input image to $352\times 352$ and the batch size to 16. The initial learning rate is 2e-4 and will be halved if the AUC on the validation set fails to increase for five epochs until the convergence. In the DFPL module, the channels of all-level features are reduced to $64$. In addition, we set the $\lambda_{r}$ and $\lambda_{e}$ for the region and edge loss to 5 and 1, respectively.
@@ IV-B Quantitative Comparisons
In-domain Localization. We first compare our DeFI-Net with several open-source methods, including MT-Net [45], HP-FCN [17], IID-Net [18], PS-Net [20], PSCC-Net [10], MVSS-Net [14]. For fairness, we train our method and all baselines using the same settings on the GC-74K and DE-24K datasets, respectively. Table II presents the localization results under in-domain testing. Our method achieves the best performance on both datasets. Specifically, our AUC, F1, and IoU scores are $1.07\%$, $2.83\%$, and $3.76\%$ higher than those of the second-place method, respectively. On the DE-3K dataset, single-scale encoding-decoding methods (such as IID-Net) experienced a notable performance drop due to the challenges in identifying small inpainting regions in this dataset. Our DeFI-Net outperforms the comparison methods significantly, particularly achieving a 23.3% lead over the second-place PSCC-Net in terms of the IoU metric. This demonstrates the effectiveness and superiority of our dense feature pyramid learning design in facilitating sufficient interaction among multi-scale features, thereby enhancing the detection ability even for small inpainting targets.
TABLE II: Quantitative comparison results (%) under in-domain testing on IID-GC and DE-3K datasets.
@@
Generalized Localization. To assess the generalization capability of the proposed DeFI-Net in localizing unknown inpainting techniques and types, we conducted a comparison. This involved applying the methods trained on the GC model to detect nine unseen inpainting models in the IID-9K dataset. The results in Table III demonstrate that our method achieved the highest performance across the averaged AUC, F1 score, and IoU metrics. It is worth noting that the PSCC-Net exhibits slightly better performance than our methods in detecting image inpainting from the LB, PM, SG, and LR models, indicating its generalizability benefited from the coarse to fine-grained multilevel supervision and progressive feature fusion strategy. We also noted the latest diffusion-based detection algorithm[15] and further compare our method with baselines on more diverse image inpainting models. Table IV shows that our dense feature network exhibits strong generalization capabilities in localizing various inpainting models, including software-based and generative AI models.
@@
Robustness Evaluation. Inpainted images will inevitably undergo various post-processing operations when disseminated. Malicious editors may also use post-processing to conceal operation traces; thus, robustness is of vital significance. To evaluate the model’s robustness to post-processing, we consider four common operations: JPEG compression, adding Gaussian noise, adding salt and pepper noise perturbation, and blurring. For each post-processing method, we conduct a comprehensive evaluation of the quality of JPEG compression (from 85 to 100), the standard deviation of Gaussian noise (from 0 to 1.5), the proportion of Salt&Pepper noise (from 0 to 0.15), and the kernel size of Gaussian blur (from 1 to 7). The results obtained from employing all pre-trained models on post-processed IID-9K data are presented in Fig 10. We can observe that our DeFI-Net exhibits great robustness to noise perturbation and JPEG compression. Yet, MVSS-Net, PS-Net, and IID-Net rely on the features of the noise stream, resulting in a weak anti-interference capability. The performance of all methods deteriorates on Gaussian blurred images. This is because blurring eliminates typical features left by the inpainting in the boundary region, effectively concealing the inpainting operation. However, the AUC of DeFI-Net still outperforms the baseline by 5% to 8%.
@@
TABLE III: Quantitative comparison results (%) under cross-domain testing on IID-9K dataset. The best performance in each column is highlighted in bold, while the second highest value is underlined.
@@
TABLE IV: Quantitative comparison (F1, %) under cross-domain testing on diverse datasets. The best performance in each column is highlighted in bold, while the second highest value is underlined. * indicates that the result is referenced from the paper [15].
@@
AutoSplice
@@
IID-Net [18] 55.40 33.98 46.66 27.17 40.80
PSCC-Net [10] 61.49 26.17 58.77 31.33 44.44
MVSS-Net [14] 41.07 42.06 55.18 41.68 44.99
PS-Net [20] 46.13 44.15 49.20 33.76 43.31
DiffForensics*[15] 50.70 25.70 - - 38.20
DeFI-Net (Ours) 67.96 48.89 77.94 43.88 59.66
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 6: Robustness evaluation on postprocessed IID-9K dataset.

### v2

@@ III-B Reverse Edge Attention Enhancement
Filling content to a specific area in an image can leave detectable traces between the altered region and its surroundings. Therefore, leveraging these edge artifacts is crucial for detecting inpainting, which the mid-low-level features can guide. Nevertheless, mid-low-level features not only contain inconsistencies caused by inpainting but also a large amount of detail and structural noise, such as texture, shape, and object boundaries. Using traditional edge extraction operators [14, 58] or convolutional structures [59] will weaken the representation of inpainting artifacts. Leveraging context-rich features that include the structural information of the restoration region as a priori can effectively suppress irrelevant semantic noise and guide boundary learning.
We develop a method for detecting and enhancing edge artifacts using reverse attention called REAE, which uses the structural details in the mid-level feature $f_{3}^{3}$ as a guide, directing the attention of low-level features toward subtle boundary changes of the inpainting area. This promotes learning distinctive features to differentiate between inpainting edges and natural boundaries. As shown in Fig. 3(a), the REAE module accepts mid-level feature $f_{3}^{3}$ and low-level features $f_{1}^{3}$ and $f_{2}^{3}$ as input. Its upper branch aggregates structural information of suspicious regions through average pooling, max pooling, and $7\times 7$ convolution. Subsequently, the sigmoid output is inverted to obtain a reverse attention map emphasizing edges, expressed by Eq. (III-B). Meanwhile, the lower branch of the REAE progressively combines $f_{3}^{3}$, $f_{2}^{3}$, and $f_{1}^{3}$ for $f_{l\&m}$ through $3\times 3$ convolution, Batch Normalization (BN), and bilinear interpolation upsampling. Finally, we use Eq. (3) to fuse the feature for edge supervision and output the edge mask $M_{e}$ through a prediction head composed of multiple convolutional layers.
@@
$Me=Pred_{edge}(f_{l\&m}\otimes Att_{re}),$ (3)
@@ III-C Spatial Adaptive Feature Fusion
Due to the nature of neural networks, high-level feature maps exhibit a significantly stronger response to the target region compared to low-level features. The difference in receptive field inevitably leads the model to bias towards the side with a more robust response during the fusion process, allocating more weights to the high-level features, which can severely suppress the role of low-level features. The latter can provide detailed edge positioning and structural information, significantly impacting the detection accuracy.
@@
To unify the feature representation of the inpainting region, we design a spatially adaptive feature fusion module (SA-FF) to combine high-level features and mid-low-level features, correcting the bias brought by response intensity. Normalization can limit the response of different features to a similar magnitude, effectively eliminating the interference of response intensity. After normalization, the model can treat multi-level features fairly, making it easier to learn effective fusion weights. Specifically, we reduce the channels of $f_{l\&m}$ and $f_{h}$ to 2 through a $1\times 1$ convolution, followed by normalization. The process can be expressed as: Eq. (4) and (5).
$F_{l\&m}^{Norm}=BN(Conv_{1}(f_{l\&m})),F_{l\&m}^{Norm}\in\mathbb{R}^{H\times W\times 2}$ (4)
$F_{h}^{Norm}=BN(Conv_{1}(f_{h})),F_{h}^{Norm}\in\mathbb{R}^{H\times W\times 2}$ (5)
@@
For different spatial positions, there exists a disparity in the contribution of high-level and mid-low-level features. High-level features, rich in context, are semantically accurate but lose many spatial and geometric details. Mid-low-level features better retain these aspects, and we hope that the model will rely more on the latter in boundary areas while leveraging the information provided by the former to fill the target area. As shown in the Fig. 5, we designed a spatially adaptive weight learner to adjust the fusion weights according to positional features, fully utilizing both advantages. Specifically, we feed $f_{h}$ used for learning weights into alternating connection blocks composed of $1\times 1$ convolution, BN, and ReLU to generate position-aware fusion weights $\textbf{W}\in\mathbb{R}^{H\times W\times 4}$. Then, the final fusion output is obtained according to element-wise multiplication and channel-wise addition, as shown in Eq. (7-9). Finally, a region mask $M_{r}$ is output through a prediction head composed of convolutional layers.
W $=((\omega_{1})_{i,j},(\omega_{2})_{i,j},(\omega_{3})_{i,j},(\omega_{4})_{i,j}),$ (6)
$i\in[1,H],j\in[1,W]$
$\vspace{-0.3cm}F^{Norm}=Cat(F_{h}[0],F_{l\&m}[0],F_{h}[1],F_{l\&m}[1])$ (7)
$\vspace{-0.3cm}\mathbf{F}^{\text{Fuse}}[k]$ $=(\mathbf{W}[2k]\otimes\mathbf{F}^{\text{Norm}}[2k])\oplus(\mathbf{W}[2k+1]$ (8)
$\otimes\mathbf{F}^{\text{Norm}}[2k+1]),k\in[0,1]$
$\vspace{-0.1cm}\mathbf{F}^{\text{Fuse}}=\text{Cat}\left(\mathbf{F}^{\text{Fuse}}[0],\mathbf{F}^{\text{Fuse}}[1]\right)$ (9)
@@
where $Cat(\cdot)$ represents the ordinary concatenation along the channel dimension. $\otimes$ and $\oplus$ denote the element-wise multiplication and element-wise addition, respectively. []delimited-[][$\cdot$][ ⋅ ] is defined as the slicing operation of the channel dimension. Note that in our design, W will learn different weights for each spatial position of each channel, totaling $4\times H\times W$ weights.
@@ III-D Loss Function
To train our DeFI-Net, we use an objective function that combines the region loss and the edge loss. The region loss enhances the model’s sensitivity to the inpainting region globally, promoting the overlap between the predicted mask and the ground truth. In contrast, the edge loss guides the model to learn non-semantic detail differences from a local perspective, achieving more precise localization. According to [63], we define the region loss function as $L_{region}=L_{bce}^{w}+L_{iou}^{w}$. Compared with the traditional Binary Cross-Entropy (BCE) loss, $L_{bce}^{w}$ assigns different weights to pixels based on their positions, paying more attention to the target area and hard pixels (e.g., boundaries prone to classification errors). The IoU loss $L_{iou}^{w}$ assigns greater weight to the target area, encouraging the network to focus on the global structure. For the edge loss, we use $L_{bce}^{w}$, similar to supervise the edge prediction, denoted as $L_{edge}$. The overall loss function is a convex combination of two losses: $L_{total}=\lambda_{r}L_{region}+\lambda_{e}L_{edge}$.
@@ IV-A Experimental Settings
Datasets. To evaluate the performance of our DeFI-Net, we use seven diverse image inpainting datasets, considering variations in inpainting model, region size, and image content. Table II summarizes the datasets. IID-84K [18] comprises 74K inpainted images created using the Gated Convolution (GC) model [39] on images sourced from Places [64] (with JPEG lossy compression) and Dresden [65] (with NEF lossless compression). Additionally, it includes a subset with 1K GC images for in-domain testing (denoted as IID-GC) and another 9K images (denoted as IID-9K) generated by four other deep learning inpainting models (CA [40], SH [34], LB [41], RN [42]) and five traditional inpainting models (TE [28], NS [2], PM [29], SG [30], LR [31]) on images from CelebA [66] and ImageNet [67]. DEFACTO [68] contains different manipulation types while our primary focus lies on 20K inpainting images and 4K authentic images (DE-24K) for training, along with a validation set (DE-1K) and a test set (DE-3K). In addition, we consider diverse datasets created by modern photo-editing software (e.g., Korus [69], Photoshop [20]) and AIGC technology (e.g., AutoSplice [46], CocoGlide [27], TGIF [70]) to evaluate the generalization ability of our method.
TABLE II: Details of image inpainting datasets in our experiment.
@@
CocoGlide [27] - - 512 GLIDE MSCOCO [72]
TGIF [70] - - 4.1K SD2, SDXL, Adobe Firefly MSCOCO [72]
@@
Implementation Details. We use the HRNet [21] initialized with pre-trained weights from ImageNet [67] as an example. Our model is implemented in Pytorch and trained using the Adam optimizer. During the training phase, we set the size of the input image to $352\times 352$ with standard data augmentation techniques, including random flipping, cropping, and rotation. The batch size is 16. The initial learning rate is 2e-4 and will be halved if the AUC on the validation set fails to increase for five epochs until the convergence. In the DFPL module, the channels of all-level features are reduced to $64$. In addition, we set the $\lambda_{r}$ and $\lambda_{e}$ for the region and edge loss to 5 and 1, respectively.
@@ IV-B Quantitative Comparisons
In-domain Localization. We first compare our DeFI-Net with several open-source methods, including MT-Net [25], HP-FCN [17], IID-Net [18], PS-Net [20], PSCC-Net [10], MVSS-Net [14]. For fairness, we train our method and all baselines using the same settings on the IID-74K and DE-24K datasets, respectively. Table III presents the localization results under in-domain testing. Our method achieves the best performance on both datasets. Specifically, our AUC, F1, and IoU scores are $1.07\%$, $2.83\%$, and $3.76\%$ higher than those of the second-place method, respectively. On the DE-3K dataset, single-scale encoding-decoding methods (such as IID-Net) experienced a notable performance drop due to the challenges in identifying small inpainting regions in this dataset. Our DeFI-Net outperforms the comparison methods significantly, particularly achieving a 23.3% lead over the second-place PSCC-Net in terms of the IoU metric. This demonstrates the effectiveness and superiority of our dense feature pyramid learning design in facilitating sufficient interaction among multi-scale features, thereby enhancing the detection ability even for small inpainting targets.
TABLE III: Quantitative comparison results (%) under in-domain testing on IID-GC and DE-3K datasets.
@@
Generalized Localization. To assess the generalization capability of the proposed DeFI-Net in localizing unknown inpainting techniques and types, we conducted a comparison. This involved applying the methods trained on the GC model to detect nine unseen inpainting models in the IID-9K dataset. The results in Table IV demonstrate that our method achieves the highest performance across the averaged AUC, F1 score, and IoU metrics. It is worth noting that the PSCC-Net exhibits slightly better performance than our methods in detecting image inpainting from the LB, PM, SG, and LR models, indicating its generalizability benefited from the coarse to fine-grained multilevel supervision and progressive feature fusion strategy. We also note the latest diffusion-based detection algorithm [15] and further compare our method with baselines on more diverse image inpainting models. Table V shows that our dense feature network exhibits strong generalization capabilities in localizing various inpainting models, including software-based and generative AI models.
Notably, DeFI-Net shows a significant performance gap compared to CAT-Net and TruFor on the TGIF dataset. During the construction of TGIF, the generated inpainted regions were spliced back into the original image using the provided masks and an additional small gradient boundary (a few pixels wide), effectively reducing artifacts. CAT-Net and TruFor incorporate additional noise-based features to aid detection, enabling them to capture subtle noise discrepancies between the generated content and the original image. In contrast, DeFI-Net primarily relies on detail differences (e.g., texture, color, edges) and global inconsistencies to identify inpainted regions, leading to reduced performance when dealing with such specific processing. In future work, we plan to explore robust noise representations to further enhance the model’s generalization capability.
@@
Robustness Evaluation. Inpainted images will inevitably undergo various post-processing operations when disseminated. Malicious editors may also use post-processing to conceal operation traces. Thus, robustness is of vital significance. To evaluate the model’s robustness to post-processing, we consider five common operations: JPEG compression, adding Gaussian noise, adding salt and pepper noise perturbation, blurring, and image resizing. For each post-processing method, we conducted a comprehensive evaluation of the quality of JPEG compression (from 85 to 100), the standard deviation of Gaussian noise (from 0 to 1.5), the proportion of Salt&Pepper noise (from 0 to 0.15), the kernel size of Gaussian blur (from 1 to 7), and the scaling factor for image resizing (ranging from 100% to 85% of the original size). The results obtained from employing all pre-trained models on post-processed IID-9K data are presented in Fig 6. We observe that our DeFI-Net exhibits great robustness to noise perturbation, JPEG compression, and slight image resizing. Yet, MVSS-Net, PS-Net, and IID-Net rely on features from the noise stream, resulting in a weak anti-interference capability. The performance of all methods deteriorates on Gaussian blurred images. This is because blurring eliminates typical features left by the inpainting in the boundary region, effectively concealing the inpainting operation. However, the AUC of DeFI-Net still outperforms the baseline by 5% to 8%.
@@
TABLE IV: Quantitative comparison results (%) under cross-domain testing on IID-9K dataset. The best performance in each column is highlighted in bold, while the second highest value is underlined.
@@
TABLE V: Quantitative comparison (F1 %) under cross-domain testing on diverse datasets. The best performance in each column is highlighted in bold, while the second highest value is underlined. † denotes the result of using open source pre-trained weights. * indicates that the result is referenced from the paper [15].
@@
Training Data&Size
AutoSplice [46]
Korus [69]
Photoshop [20]
DE-3k [68]
CocoGlide
@@
IID-Net [18] IID-74K 55.40 33.98 46.66 27.17 56.74 40.16 43.35
PSCC-Net [10] IID-74K 61.49 26.17 58.77 31.33 52.63 36.58 44.49
MVSS-Net [14] IID-74K 41.07 42.06 55.18 41.68 50.55 39.56 45.01
PS-Net [20] IID-74K 46.13 44.15 49.20 33.76 49.18 38.16 43.43
CAT-Net[53] CASIAv2,etc -870K 68.50 48.46 64.12 50.69 55.20 64.00 58.49
TruFor[27] CASIAv2,etc -870K 72.63 48.05 47.83 42.34 62.64 50.12 53.93
DiffForensics*[15] CASIAv2,etc -49K 50.70 25.70 - - - - -
DeFI-Net (Ours) IID-74K 67.96 48.89 77.94 43.88 57.13 40.95 56.13
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 6: Robustness evaluation on postprocessed IID-9K dataset. Best viewed in color and with zoom-in.